# Invariants of persistence modules defined by order-embeddings

### v2

@@ 2.1. Definitions and first examples
A $\mathbb{Z}$-invariant (or invariant for short) on $\operatorname{\mathrm{mod}}\mathcal{P}$ is a $\mathbb{Z}$-linear map $\Phi:\mathrm{K}_{0}^{\mathrm{sp}}(\mathcal{P})\longrightarrow\mathbb{Z}^{I}$ for some set $I$. In other words, $\Phi$ can be viewed as a map $\operatorname{\mathrm{mod}}\mathcal{P}\to\mathbb{Z}^{I}$ which is constant on isoclasses of modules and such that $\Phi(M\oplus N)=\Phi(M)+\Phi(N)$. An invariant $\Phi$ is finite if its image is finitely generated. We denote by $\operatorname{\mathrm{rk}}\Phi$ the rank of its image. We say that $\Phi:\mathrm{K}_{0}^{\mathrm{sp}}(\mathcal{P})\longrightarrow\mathbb{Z}^{I}$ is a complete invariant if it is of full rank, that is, $\Phi$ is injective.
@@
    Counting the multiplicities of all indecomposable direct summands yields an invariant $\mathbf{mult}_{\mathcal{P}}:\mathrm{K}_{0}^{\mathrm{sp}}(\mathcal{P})\to\mathbb{Z}^{\operatorname{\mathrm{Ind}}\mathcal{P}}$ of rank $|\operatorname{\mathrm{Ind}}\mathcal{P}|$. It is defined by $\mathbf{mult}_{\mathcal{P}}(M)_{U}=\alpha_{U}$ where $M\simeq\bigoplus_{U\in\operatorname{\mathrm{Ind}}\mathcal{P}}U^{\alpha_{U}}$ is the decomposition of $M$ into indecomposable objects. If $\mathcal{P}$ is a poset of finite representation type, then $\mathbf{mult}_{\mathcal{P}}$ is a finite invariant. More generally, if $\mathcal{U}\subseteq\operatorname{\mathrm{Ind}}\mathcal{P}$ is a finite family of indecomposable modules, we define $\mathbf{mult}_{\mathcal{P}}^{\mathcal{U}}:\mathrm{K}_{0}^{\mathrm{sp}}(\mathcal{P})\to\mathbb{Z}^{\mathcal{U}}$ by composing $\mathbf{mult}_{\mathcal{P}}$ with the natural projection of $\mathbb{Z}^{\mathcal{U}}\to\mathbb{Z}^{\operatorname{\mathrm{Ind}}\mathcal{P}}$.
@@
    For $a\in\mathcal{P}$, $\operatorname{\mathbf{dim}}_{a}:\mathrm{K}_{0}^{\mathrm{sp}}(\mathcal{P})\to\mathbb{Z}$ is an invariant of rank $1$ given by $\operatorname{\mathbf{dim}}_{a}(M)=\dim_{k}M_{a}$. The map $\operatorname{\mathbf{dim}}_{\mathcal{P}}:=\bigoplus_{a\in\mathcal{P}}\operatorname{\mathbf{dim}}_{a}$ is an invariant of rank $|\mathcal{P}|$. This invariant is often known as the dimension vector or Hilbert function.
@@
    $\{\leq_{\mathcal{P}}\}:=\{(a,b)\in\mathcal{P}^{2}\mid a\leq b\}.$
    For $(a,b)\in\{\leq_{\mathcal{P}}\}$, the map $\operatorname{\mathbf{rk}}_{a,b}:\mathrm{K}_{0}^{\mathrm{sp}}(\mathcal{P})\to\mathbb{Z}$ given by $\operatorname{\mathbf{rk}}_{a,b}(M)=\operatorname{\mathrm{rk}}M_{a\to b}$ is an invariant of rank $1$. The map $\operatorname{\mathbf{rk}}_{\mathcal{P}}:=\bigoplus_{(a,b)\in\{\leq_{\mathcal{P}}\}}\operatorname{\mathbf{rk}}_{a,b}$ is an invariant of rank |$|\{\leq_{\mathcal{P}}\}|$ called the rank invariant (see [CZ09, BOO24]).
@@
    For any finite family $\mathcal{U}\subseteq\operatorname{\mathrm{Ind}}\mathcal{P}$, the map $\mathbf{dimh}_{\mathcal{P}}^{\mathcal{U}}:\mathrm{K}_{0}^{\mathrm{sp}}(\mathcal{P})\to\mathbb{Z}^{\mathcal{U}}$ defined by $\left(\mathbf{dimh}_{\mathcal{P}}^{\mathcal{U}}(M)\right)_{U}:=\dim_{k}\operatorname{\mathrm{Hom}}_{\mathcal{P}}(U,M)$ for $U\in\mathcal{U}$ is an invariant of rank $|\mathcal{U}|$. Following [BBH24], invariants of this form are called dim-Hom invariants.
@@ 2.2. Comparing invariants
Two invariants $\Phi_{1},\Phi_{2}$ on $\operatorname{\mathrm{mod}}\mathcal{P}$ are equivalent (in symbols $\Phi_{1}\simeq\Phi_{2}$) if there exists an isomorphism $\varphi:\operatorname{\mathrm{Im}}\Phi_{1}\to\operatorname{\mathrm{Im}}\Phi_{2}$ such that $\varphi\circ\Phi_{1}=\Phi_{2}$.
@@
Let $\Phi_{1},\Phi_{2}$ be two invariants on $\operatorname{\mathrm{mod}}\mathcal{P}$. We say that $\Phi_{1}$ is finer than $\Phi_{2}$, and denote it by $\Phi_{1}\geq\Phi_{2}$, if there exists a $\mathbb{Z}$-linear map $\varphi:\operatorname{\mathrm{Im}}\Phi_{1}\to\operatorname{\mathrm{Im}}\Phi_{2}$ such that $\varphi\circ\Phi_{1}=\Phi_{2}$ (in particular, $\varphi$ is surjective). This defines a partial order on the equivalence classes of invariants on $\operatorname{\mathrm{mod}}\mathcal{P}$.
@@
    The diagonal inclusion $\mathcal{P}\to\{\leq_{\mathcal{P}}\}$ induces an injective map $\varphi:\mathbb{Z}^{\mathcal{P}}\to\mathbb{Z}^{\{\leq_{\mathcal{P}}\}}$ whose transpose $\varphi^{T}:\mathbb{Z}^{\{\leq_{\mathcal{P}}\}}\to\mathbb{Z}^{\mathcal{P}}$ is surjective and satisfies $\varphi^{T}\circ\operatorname{\mathbf{rk}}_{\mathcal{P}}=\operatorname{\mathbf{dim}}_{\mathcal{P}}$, hence $\operatorname{\mathbf{rk}}_{\mathcal{P}}\geq\operatorname{\mathbf{dim}}_{\mathcal{P}}$. In general, the rank invariant is strictly finer than the dimension vector, in fact for a connected poset $\mathcal{P}$ the rank invariant is equivalent to the dimension vector if and only if $|\mathcal{P}|=1$. It is a classical result in representation theory that the rank invariant $\operatorname{\mathbf{rk}}_{\mathcal{P}}$ is complete for a totally ordered poset $\mathcal{P}$.
@@
Denote by $\{U_{1},\ldots,U_{\ell}\}=\operatorname{\mathrm{Ind}}\mathcal{P}$. For $1\leq i,j\leq\ell$ we have $\mathbf{mult}_{U_{i}}(U_{j})=\delta_{i,j}$ and $\mathbf{dimh}_{U_{i}}(U_{j})=\dim_{k}\operatorname{\mathrm{Hom}}_{\mathcal{P}}(U_{i},U_{j})$. Hence if we denote by $\Phi\in{\rm Mat}_{\ell}(\mathbb{Z})$ the matrix defined by $\Phi_{i,j}:=\dim_{k}\operatorname{\mathrm{Hom}}_{\mathcal{P}}(U_{i},U_{j})$, we have that $\Phi\circ\mathbf{mult}_{\mathcal{P}}=\mathbf{dimh}_{\mathcal{P}}^{\operatorname{\mathrm{Ind}}\mathcal{P}}$ showing that $\mathbf{mult}_{\mathcal{P}}\geq\mathbf{dimh}_{\mathcal{P}}^{\operatorname{\mathrm{Ind}}\mathcal{P}}$.
Since $\mathcal{P}$ is representation-finite, it is representation-directed by [Si10, Theorem 1.4]. This means that the transitive closure of the relation $U_{i}\leq U_{j}$ whenever $\operatorname{\mathrm{Hom}}(U_{i},U_{j})\neq 0$ is a partial order on $\operatorname{\mathrm{Ind}}\mathcal{P}$. It also implies that every indecomposable module is a brick; that is, that $\operatorname{\mathrm{End}}_{\mathcal{P}}(U_{i})\cong k$ for all $U_{i}\in\operatorname{\mathrm{Ind}}\mathcal{P}$. We conclude that $\Phi$ is an isomorphism using the following lemma with $Y=\operatorname{\mathrm{Ind}}\mathcal{P}$ and $\varphi_{i}(j)=\Phi_{i,j}.$
@@
The proof is classical, but since we use this result in several places in the paper, we give here a proof for the convenience of the reader. We first show that the family $(\varphi_{y})_{y\in Y}$ generates the free abelian group $\mathbb{Z}^{Y}$. For $\Psi\in\mathbb{Z}^{Y}$, denote
$\mathcal{S}_{\Psi}:=\{y\in Y\mid\exists y^{\prime}\geq y\textrm{ with }\Psi(y^{\prime})\neq 0\}.$
We show by induction on $|\mathcal{S}_{\Psi}|$ that $\Psi$ is a linear combination of the $\varphi_{y}$ with $y\in\mathcal{S}_{\Psi}$. If $\mathcal{S}_{\Psi}=\emptyset$ then $\Psi=0$ and there is nothing to show. Assume now that $\mathcal{S}_{\Psi}\neq\emptyset$ and let $y\in\mathcal{S}_{\Psi}$ be a maximal element, hence $\Psi(y)\neq 0.$ Define $\Psi^{\prime}:=\Psi-\Psi(y)\varphi_{y}$. We will show that $\mathcal{S}_{\Psi^{\prime}}\subseteq\mathcal{S}_{\Psi}$. Let $z$ be in $\mathcal{S}_{\Psi^{\prime}}$. Then there exists $y^{\prime}\geq z$ such that $\Psi^{\prime}(y^{\prime})\neq 0$. That is $\Psi(y^{\prime})\neq\Psi(y)\varphi_{y}(y^{\prime})$. If $\Psi(y^{\prime})\neq 0$ then $z\in\mathcal{S}_{\Psi}$. If $\Psi(y^{\prime})=0$, then $\varphi_{y}(y^{\prime})\neq 0$ which implies by hypothesis that $y\geq y^{\prime}$. Hence we have $y\geq z$ with $\Psi(y)\neq 0$, that is $z\in\mathcal{S}_{\Psi}$. Hence we have $\mathcal{S}_{\Psi^{\prime}}\subseteq\mathcal{S}_{\Psi}$. Moreover $y\in\mathcal{S}_{\Psi}$ but $y\notin\mathcal{S}_{\Psi^{\prime}}$. Thus we know by induction that $\Psi^{\prime}$ is a linear combination of the $\varphi_{y}$ with $y\in\mathcal{S}_{\Psi}^{\prime}$, and so $\Psi=\Psi^{\prime}+\Psi(y)\varphi_{y}$ is a linear combination of the $\varphi_{y}$ with $y\in\mathcal{S}_{\Psi}$.
Let $\sum_{y\in Y}\alpha_{y}\varphi_{y}=0$ be a linear combination. Assume that $\{y\in Y\mid\alpha_{y}\neq 0\}$ is non empty and let $z$ be a maximal element in this set. Then we obtain $-\alpha_{z}=\sum_{y\neq z}\alpha_{y}\varphi_{y}(z)\neq 0$. Therefore there exists $y$ with $\alpha_{y}\neq 0$ and $\varphi_{y}(z)\neq 0$ which implies $z\leq y$ and contradicts the maximality of $z$. Hence the family is linearly independent. ∎
@@ 3. Invariants defined by order-embeddings
Let $\mathcal{X}$ and $\mathcal{P}$ be finite posets. A map $f:\mathcal{X}\to\mathcal{P}$ is an order-embedding if $x\leq x^{\prime}$ precisely when $f(x)\leq f(x^{\prime})$, for any $x,x^{\prime}\in\mathcal{X}$. Denote by $\mathrm{Emb}_{\mathcal{X}}^{\mathcal{P}}$ the set of all order-embeddings of $\mathcal{X}$ into $\mathcal{P}$ up to automorphism of $\mathcal{X}$.
@@
Any $f\in\mathrm{Emb}_{\mathcal{X}}^{\mathcal{P}}$ induces a restriction functor $f^{*}:\operatorname{\mathrm{mod}}\mathcal{P}\to\operatorname{\mathrm{mod}}\mathcal{X}$ which is exact. It is also dense since order-embeddings are injective. Therefore $f^{*}$ induces a surjective $\mathbb{Z}$-linear map $\mathrm{K}_{0}^{\mathrm{sp}}(\mathcal{P})\to\mathrm{K}_{0}^{\mathrm{sp}}(\mathcal{X})$.
@@
Let $\mathcal{X}$ be a poset of finite representation type, $\mathcal{P}$ be a finite poset and, $\mathrm{E}$ be a subset of $\mathrm{Emb}_{\mathcal{X}}^{\mathcal{P}}$. We denote by $\mathrm{I}:=\operatorname{\mathrm{Ind}}\mathcal{X}$.
We define $\mathbf{mult}_{\mathcal{X},\mathcal{P}}^{\mathrm{E}}:\mathrm{K}_{0}^{\mathrm{sp}}(\mathcal{P})\to\mathbb{Z}^{\mathrm{E}\times\mathrm{I}}$ by
$\left(\mathbf{mult}_{\mathcal{X},\mathcal{P}}^{\mathrm{E}}(M)\right)_{(f,U)}:={\rm mult}_{\mathcal{X}}(U,f^{*}M)\quad\textrm{for }(f,U)\in\mathrm{E}\times\mathrm{I}.$
We define $\mathbf{dimh}^{E}_{\mathcal{X},\mathcal{P}}:\mathrm{K}_{0}^{\mathrm{sp}}(\mathcal{P})\to\mathbb{Z}^{E\times I}$ by
$\left(\mathbf{dimh}_{\mathcal{X},\mathcal{P}}^{E}(M)\right)_{(f,U)}:=\dim\operatorname{\mathrm{Hom}}_{\mathcal{X}}(U,f^{*}M)\quad\textrm{for }(f,U)\in\mathrm{E}\times\mathrm{I}.$
@@
    $\left(\mathbf{mult}_{\mathcal{X},\mathcal{P}}^{\mathrm{Emb}_{\mathcal{X}}^{\mathcal{P}}}(M)\right)_{(f,U)}=0\text{ for all $(f,U)\in\mathrm{Emb}_{\mathcal{X}}^{\mathcal{P}}\times\{\mathbb{I}_{\{a,b,c\}},\mathbb{I}_{\{b,c,d\}}\}$}.$
@@
By Proposition 2.7, there is an invertible map $\Phi:\mathbb{Z}^{\mathrm{I}}\to\mathbb{Z}^{\mathrm{I}}$ such that $\Phi\circ\mathbf{mult}_{\mathcal{X}}=\mathbf{dimh}_{\mathcal{X}}^{\mathrm{I}}$. We then obtain the following commutative diagram
[…]
where $\Phi^{\mathrm{E}}$ is an isomorphism. ∎
@@
    Let $f:\mathcal{X}_{2}\to\mathcal{P}$ be an order-embedding. Then one easily checks that for any $M\in\operatorname{\mathrm{mod}}\mathcal{P}$
    $\mathbf{mult}_{\mathcal{X}_{2},\mathcal{P}}(M)_{(f,U)}=\left\{\begin{array}[]{ll}\dim M_{f2}-\operatorname{\mathrm{rk}}M_{f1\to f2}&\textrm{ if }U=\mathbb{I}_{\{2\}}\\ \operatorname{\mathrm{rk}}M_{f1\to f2}&\textrm{ if }U=\mathbb{I}_{\{1,2\}}\\ \dim M_{f1}-\operatorname{\mathrm{rk}}M_{f1\to f2}&\textrm{ if }U=\mathbb{I}_{\{1\}}\end{array}\right.$
@@
    $\mathbf{dimh}_{\mathcal{X}_{2},\mathcal{P}}(M)_{(f,U)}=\left\{\begin{array}[]{ll}\dim M_{f2}&\textrm{ if }U=\mathbb{I}_{\{2\}}\\ \dim M_{f1}&\textrm{ if }U=\mathbb{I}_{\{1,2\}}\\ \dim M_{f1}-\operatorname{\mathrm{rk}}M_{f1\to f2}&\textrm{ if }U=\mathbb{I}_{\{1\}}\end{array}\right.$
@@
    $\Phi=\begin{pmatrix}1&1&0\\ 0&1&1\\ 0&0&1\end{pmatrix}.$
From this example we deduce the following

### v3

@@ 2.1. Definitions and first examples
A $\mathbb{Z}$-invariant (or invariant for short) on $\operatorname{\mathrm{mod}}\mathcal{P}$ is a $\mathbb{Z}$-linear map $\Phi\mathrel{\mathop{\ordinarycolon}}\mathrm{K}_{0}^{\mathrm{sp}}(\mathcal{P})\longrightarrow\mathbb{Z}^{I}$ for some set $I$. In other words, $\Phi$ can be viewed as a map $\operatorname{\mathrm{mod}}\mathcal{P}\to\mathbb{Z}^{I}$ which is constant on isoclasses of modules and such that $\Phi(M\oplus N)=\Phi(M)+\Phi(N)$. An invariant $\Phi$ is finite if its image is finitely generated. We denote by $\operatorname{\mathrm{rk}}\Phi$ the rank of its image. We say that $\Phi\mathrel{\mathop{\ordinarycolon}}\mathrm{K}_{0}^{\mathrm{sp}}(\mathcal{P})\longrightarrow\mathbb{Z}^{I}$ is a complete invariant if $\Phi$ is injective.
@@
    Counting the multiplicities of all indecomposable direct summands yields an invariant $\mathbf{mult}_{\mathcal{P}}\mathrel{\mathop{\ordinarycolon}}\mathrm{K}_{0}^{\mathrm{sp}}(\mathcal{P})\to\mathbb{Z}^{\operatorname{\mathrm{Ind}}\mathcal{P}}$ of rank $|\operatorname{\mathrm{Ind}}\mathcal{P}|$. It is defined by $\mathbf{mult}_{\mathcal{P}}(M)_{U}=\alpha_{U}$ where $M\simeq\bigoplus_{U\in\operatorname{\mathrm{Ind}}\mathcal{P}}U^{\alpha_{U}}$ is the decomposition of $M$ into indecomposable objects. If $\mathcal{P}$ is a poset of finite representation type, then $\mathbf{mult}_{\mathcal{P}}$ is a finite invariant. More generally, if $\mathcal{U}\subseteq\operatorname{\mathrm{Ind}}\mathcal{P}$ is a finite family of indecomposable modules, we define $\mathbf{mult}_{\mathcal{P}}^{\mathcal{U}}\mathrel{\mathop{\ordinarycolon}}\mathrm{K}_{0}^{\mathrm{sp}}(\mathcal{P})\to\mathbb{Z}^{\mathcal{U}}$ by composing $\mathbf{mult}_{\mathcal{P}}$ with the natural projection of $\mathbb{Z}^{\mathcal{U}}\to\mathbb{Z}^{\operatorname{\mathrm{Ind}}\mathcal{P}}$.
@@
    For $a\in\mathcal{P}$, $\operatorname{\mathbf{dim}}_{a}\mathrel{\mathop{\ordinarycolon}}\mathrm{K}_{0}^{\mathrm{sp}}(\mathcal{P})\to\mathbb{Z}$ is an invariant of rank $1$ given by $\operatorname{\mathbf{dim}}_{a}(M)=\dim_{k}M_{a}$. The map $\operatorname{\mathbf{dim}}_{\mathcal{P}}\mathrel{\mathop{\ordinarycolon}}=\bigoplus_{a\in\mathcal{P}}\operatorname{\mathbf{dim}}_{a}$ is an invariant of rank $|\mathcal{P}|$. This invariant is often known as the dimension vector or Hilbert function.
@@
    $\{\leq_{\mathcal{P}}\}\mathrel{\mathop{\ordinarycolon}}=\{(a,b)\in\mathcal{P}^{2}\mid a\leq b\}.$
    For $(a,b)\in\{\leq_{\mathcal{P}}\}$, the map $\operatorname{\mathbf{rk}}_{a,b}\mathrel{\mathop{\ordinarycolon}}\mathrm{K}_{0}^{\mathrm{sp}}(\mathcal{P})\to\mathbb{Z}$ given by $\operatorname{\mathbf{rk}}_{a,b}(M)=\operatorname{\mathrm{rk}}M_{a\to b}$ is an invariant of rank $1$. The map $\operatorname{\mathbf{rk}}_{\mathcal{P}}\mathrel{\mathop{\ordinarycolon}}=\bigoplus_{(a,b)\in\{\leq_{\mathcal{P}}\}}\operatorname{\mathbf{rk}}_{a,b}$ is an invariant of rank $|\{\leq_{\mathcal{P}}\}|$ called the rank invariant (see [CZ09, BOO24]).
@@
    For any finite family $\mathcal{U}\subseteq\operatorname{\mathrm{Ind}}\mathcal{P}$, the map $\mathbf{dimh}_{\mathcal{P}}^{\mathcal{U}}\mathrel{\mathop{\ordinarycolon}}\mathrm{K}_{0}^{\mathrm{sp}}(\mathcal{P})\to\mathbb{Z}^{\mathcal{U}}$ defined by $\left(\mathbf{dimh}_{\mathcal{P}}^{\mathcal{U}}(M)\right)_{U}\mathrel{\mathop{\ordinarycolon}}=\dim_{k}\operatorname{\mathrm{Hom}}_{\mathcal{P}}(U,M)$ for $U\in\mathcal{U}$ is an invariant of rank $|\mathcal{U}|$. Following [BBH24], invariants of this form are called dim-Hom invariants.
@@ 2.2. Comparing invariants
Two invariants $\Phi_{1},\Phi_{2}$ on $\operatorname{\mathrm{mod}}\mathcal{P}$ are equivalent (in symbols $\Phi_{1}\simeq\Phi_{2}$) if there exists an isomorphism $\varphi\mathrel{\mathop{\ordinarycolon}}\operatorname{\mathrm{Im}}\Phi_{1}\to\operatorname{\mathrm{Im}}\Phi_{2}$ such that $\varphi\circ\Phi_{1}=\Phi_{2}$.
@@
Let $\Phi_{1},\Phi_{2}$ be two invariants on $\operatorname{\mathrm{mod}}\mathcal{P}$. We say that $\Phi_{1}$ is finer than $\Phi_{2}$, and denote it by $\Phi_{1}\geq\Phi_{2}$, if there exists a $\mathbb{Z}$-linear map $\varphi\mathrel{\mathop{\ordinarycolon}}\operatorname{\mathrm{Im}}\Phi_{1}\to\operatorname{\mathrm{Im}}\Phi_{2}$ such that $\varphi\circ\Phi_{1}=\Phi_{2}$ (in particular, $\varphi$ is surjective). This defines a partial order on the equivalence classes of invariants on $\operatorname{\mathrm{mod}}\mathcal{P}$.
@@
    The diagonal inclusion $\mathcal{P}\to\{\leq_{\mathcal{P}}\}$ induces an injective map $\varphi\mathrel{\mathop{\ordinarycolon}}\mathbb{Z}^{\mathcal{P}}\to\mathbb{Z}^{\{\leq_{\mathcal{P}}\}}$ whose transpose $\varphi^{T}\mathrel{\mathop{\ordinarycolon}}\mathbb{Z}^{\{\leq_{\mathcal{P}}\}}\to\mathbb{Z}^{\mathcal{P}}$ is surjective and satisfies $\varphi^{T}\circ\operatorname{\mathbf{rk}}_{\mathcal{P}}=\operatorname{\mathbf{dim}}_{\mathcal{P}}$, hence $\operatorname{\mathbf{rk}}_{\mathcal{P}}\geq\operatorname{\mathbf{dim}}_{\mathcal{P}}$. In general, the rank invariant is strictly finer than the dimension vector, in fact for a connected poset $\mathcal{P}$ the rank invariant is equivalent to the dimension vector if and only if $|\mathcal{P}|=1$. It is a classical result in representation theory that the rank invariant $\operatorname{\mathbf{rk}}_{\mathcal{P}}$ is complete for a totally ordered poset $\mathcal{P}$.
@@
Denote by $\{U_{1},\ldots,U_{\ell}\}=\operatorname{\mathrm{Ind}}\mathcal{P}$. For $1\leq i,j\leq\ell$ we have $\mathbf{mult}_{U_{i}}(U_{j})=\delta_{i,j}$ and $\mathbf{dimh}_{U_{i}}(U_{j})=\dim_{k}\operatorname{\mathrm{Hom}}_{\mathcal{P}}(U_{i},U_{j})$. Hence if we denote by $\Phi\in{\rm Mat}_{\ell}(\mathbb{Z})$ the matrix defined by $\Phi_{i,j}\mathrel{\mathop{\ordinarycolon}}=\dim_{k}\operatorname{\mathrm{Hom}}_{\mathcal{P}}(U_{i},U_{j})$, we have that $\Phi\circ\mathbf{mult}_{\mathcal{P}}=\mathbf{dimh}_{\mathcal{P}}^{\operatorname{\mathrm{Ind}}\mathcal{P}}$ showing that $\mathbf{mult}_{\mathcal{P}}\geq\mathbf{dimh}_{\mathcal{P}}^{\operatorname{\mathrm{Ind}}\mathcal{P}}$.
Since $\mathcal{P}$ is representation-finite, it is representation-directed by [Si10, Theorem 1.4]. This means that the transitive closure of the relation $U_{i}\leq U_{j}$ whenever $\operatorname{\mathrm{Hom}}(U_{i},U_{j})\neq 0$ is a partial order on $\operatorname{\mathrm{Ind}}\mathcal{P}$. This in turn implies that every indecomposable module admits only the trivial endomorphism, thus it is a brick. We conclude that $\Phi$ is an isomorphism using the following lemma with $Y=\operatorname{\mathrm{Ind}}\mathcal{P}$ and $\varphi_{i}(j)=\Phi_{i,j}.$
@@
Label the elements $\{y_{1},\ldots,y_{n}\}$ of $Y$ in such a way that $y_{i}\leq y_{j}$ implies $i\leq j$. Consider the dual basis $(\delta_{y_{i}})_{i=1,\ldots,n}$ of $\mathbb{Z}^{Y}$ defined by $\delta_{y_{i}}(y_{j})=\delta_{i,j}$. Then the matrix expressing the $(\phi_{y_{i}})_{i=1,\ldots,n}$ in the basis $(\delta_{y_{i}})_{i=1,\ldots,n}$ is in ${\rm GL}_{n}(\mathbb{Z})$ since it is an upper triangular matrix with $1$ on the diagonal. Hence the family $(\varphi_{y_{i}})_{i=1,\ldots,n}$ is a $\mathbb{Z}$-basis. ∎
@@ 3. Invariants defined by order-embeddings
Let $\mathcal{X}$ and $\mathcal{P}$ be finite posets. A map $f\mathrel{\mathop{\ordinarycolon}}\mathcal{X}\to\mathcal{P}$ is an order-embedding if $x\leq x^{\prime}$ precisely when $f(x)\leq f(x^{\prime})$, for any $x,x^{\prime}\in\mathcal{X}$. Denote by $\mathrm{Emb}_{\mathcal{X}}^{\mathcal{P}}$ the set of all order-embeddings of $\mathcal{X}$ into $\mathcal{P}$ up to automorphism of $\mathcal{X}$.
@@
Any $f\in\mathrm{Emb}_{\mathcal{X}}^{\mathcal{P}}$ induces a restriction functor $f^{*}\mathrel{\mathop{\ordinarycolon}}\operatorname{\mathrm{mod}}\mathcal{P}\to\operatorname{\mathrm{mod}}\mathcal{X}$ which is exact. It is also dense since order-embeddings are injective. Therefore $f^{*}$ induces a surjective $\mathbb{Z}$-linear map $\mathrm{K}_{0}^{\mathrm{sp}}(\mathcal{P})\to\mathrm{K}_{0}^{\mathrm{sp}}(\mathcal{X})$.
@@
Let $\mathcal{X}$ be a poset of finite representation type, $\mathcal{P}$ be a finite poset and, $\mathrm{E}$ be a subset of $\mathrm{Emb}_{\mathcal{X}}^{\mathcal{P}}$. We denote by $\mathrm{I}\mathrel{\mathop{\ordinarycolon}}=\operatorname{\mathrm{Ind}}\mathcal{X}$.
We define $\mathbf{mult}_{\mathcal{X},\mathcal{P}}^{\mathrm{E}}\mathrel{\mathop{\ordinarycolon}}\mathrm{K}_{0}^{\mathrm{sp}}(\mathcal{P})\to\mathbb{Z}^{\mathrm{E}\times\mathrm{I}}$ by
$\left(\mathbf{mult}_{\mathcal{X},\mathcal{P}}^{\mathrm{E}}(M)\right)_{(f,U)}\mathrel{\mathop{\ordinarycolon}}={\rm mult}_{\mathcal{X}}(U,f^{*}M)\quad\textrm{for }(f,U)\in\mathrm{E}\times\mathrm{I}.$
We define $\mathbf{dimh}^{E}_{\mathcal{X},\mathcal{P}}\mathrel{\mathop{\ordinarycolon}}\mathrm{K}_{0}^{\mathrm{sp}}(\mathcal{P})\to\mathbb{Z}^{E\times I}$ by
$\left(\mathbf{dimh}_{\mathcal{X},\mathcal{P}}^{E}(M)\right)_{(f,U)}\mathrel{\mathop{\ordinarycolon}}=\dim\operatorname{\mathrm{Hom}}_{\mathcal{X}}(U,f^{*}M)\quad\textrm{for }(f,U)\in\mathrm{E}\times\mathrm{I}.$
@@
    $\left(\mathbf{mult}_{\mathcal{X},\mathcal{P}}(M)\right)_{(f,U)}=0\text{ for all $(f,U)\in\mathrm{Emb}_{\mathcal{X}}^{\mathcal{P}}\times\{\mathbb{I}_{\{a,b,c\}},\mathbb{I}_{\{b,c,d\}}\}$}.$
@@
By Proposition 2.7, there is an invertible map $\Phi\mathrel{\mathop{\ordinarycolon}}\mathbb{Z}^{\mathrm{I}}\to\mathbb{Z}^{\mathrm{I}}$ such that $\Phi\circ\mathbf{mult}_{\mathcal{X}}=\mathbf{dimh}_{\mathcal{X}}^{\mathrm{I}}$. We then obtain the following commutative diagram
${\mathrm{K}_{0}^{\mathrm{sp}}(\mathcal{P})\ignorespaces\ignorespaces\ignorespaces\ignorespaces\ignorespaces\ignorespaces\ignorespaces\ignorespaces}$$\scriptstyle{\mathbf{mult}_{\mathcal{X},\mathcal{P}}^{\mathrm{E}}}$${\mathbb{Z}^{\mathrm{E}\times I}\ignorespaces\ignorespaces\ignorespaces\ignorespaces}$$\scriptstyle{\Phi^{\mathrm{E}}}$${\mathrm{K}_{0}^{\mathrm{sp}}(\mathcal{P})\ignorespaces\ignorespaces\ignorespaces\ignorespaces}$$\scriptstyle{\mathbf{dimh}_{\mathcal{X},\mathcal{P}}^{\mathrm{E}}}$${\mathbb{Z}^{\mathrm{E}\times\mathrm{I}}}$
where $\Phi^{\mathrm{E}}$ is the isomorphism sending $x\in\mathbb{Z}^{\mathrm{E}\times\mathrm{I}}$ to the map $\Phi^{\mathrm{E}}(x)\in\mathbb{Z}^{\mathrm{E}\times\mathrm{I}}$ which is defined by applying, for a fixed $f\in\mathrm{E}$, the isomorphism $\Phi$ to the map $x_{(f,-)}\in\mathbb{Z}^{\mathrm{I}}$; in formulas:
$\Phi^{\mathrm{E}}(x)_{(f,M)}\mathrel{\mathop{\ordinarycolon}}=\Phi(x_{(f,-)})_{M}\quad\textrm{for }f\in\mathrm{E},\ M\in\mathrm{I},\ x\in\mathbb{Z}^{\mathrm{E}\times\mathrm{I}}.$
@@
    Let $f\mathrel{\mathop{\ordinarycolon}}\mathcal{X}_{2}\to\mathcal{P}$ be an order-embedding. Then one easily checks that for any $M\in\operatorname{\mathrm{mod}}\mathcal{P}$
    $\mathbf{mult}_{\mathcal{X}_{2},\mathcal{P}}(M)_{(f,U)}=\left\{\begin{array}[]{ll}\dim M_{f2}-\operatorname{\mathrm{rk}}M_{f1\to f2}&\textrm{ if }U=\mathbb{I}_{\{2\}}\\ \operatorname{\mathrm{rk}}M_{f1\to f2}&\textrm{ if }U=\mathbb{I}_{\{1,2\}}\\ \dim M_{f1}-\operatorname{\mathrm{rk}}M_{f1\to f2}&\textrm{ if }U=\mathbb{I}_{\{1\}}\end{array}\right.\qquad(3.7.1)$
@@
    $\mathbf{dimh}_{\mathcal{X}_{2},\mathcal{P}}(M)_{(f,U)}=\left\{\begin{array}[]{ll}\dim M_{f2}&\textrm{ if }U=\mathbb{I}_{\{2\}}\\ \dim M_{f1}&\textrm{ if }U=\mathbb{I}_{\{1,2\}}\\ \dim M_{f1}-\operatorname{\mathrm{rk}}M_{f1\to f2}&\textrm{ if }U=\mathbb{I}_{\{1\}}\end{array}\right.\qquad(3.7.2)$
@@
    $\Phi=\begin{pmatrix}1&0&0\\ 1&1&0\\ 0&1&1\end{pmatrix}$.
From part (2) of this example we can also deduce that multiplicities can be calculated from ranks: Equation (3.7.1) shows that composing the vector of ranks with the matrix
$\Phi_{f}=\begin{pmatrix}1&-1&0\\ 0&\;\;\;1&0\\ 0&-1&1\end{pmatrix}$
yields the vector of multiplicities.
We formalize this observation in the following corollary: